# Dual-component stellar assembly histories in local elliptical galaxies via MUSE

### v1

@@ 4.2 Mean stellar populations
Figure 4 shows the stellar populations obtained from full spectral fitting with pPXF (see Sect. 3.2). The mean mass-weighted stellar metallicities [M/H] and the logarithmic stellar ages, are depicted in the upper panel, while the mean light-weighted properties are shown in the lower panel. The circles, squares, and triangles represent the three elliptical galaxies in the sample (J020536, J205050, and J225546 respectively), and the red and blue colours mark the inner and outer components derived from BUDDI in the two-component (Sérsic + Sérsic) models. The inner and outer components of the same galaxy are connected by the black lines for better readability. For comparison, the stellar population properties from the standard single component (Sérsic) model are shown in purple. Mass-weighted ages reflect the period when the majority of stellar mass was assembled and are less sensitive to the most luminous and younger stars; whereas light-weighted ages are influenced by recent star formation, which significantly impacts luminosity but contributes less to the overall mass (Hopkins, 2018).
@@
The mean mass and light-weighted ages of the galaxies (single Sérsic) and their individual components (inner and outer components in the double Sérsic model) are all older than 9 Gyr, which typically places them in the category of long-term quiescent “classical ellipticals”. The difference between the ages of the populations present in the inner and outer components ranges roughly between 1 and 5 Gyr. Plauchu-Frayn et al. (2012) define the star formation timescale as the difference between the mass-weighted and light-weighted stellar ages: Gyr. is high for objects that are continuously forming stars, while it is low for those that have ceased forming stars a long time ago. Following their definition, we compute for each component, as well as globally for each galaxy. Considering the ages of the galaxies as a whole, we find to be negligible, of the order of a few 100 Myr, well within their uncertainties. This insignificant difference is typically expected for classical ellipticals when considered as a global property, implying that their stellar populations are homogeneous in age and would have formed and quenched at roughly similar early lookback times. The inner and outer components similarly show virtually imperceptible values. While we recognise that the mean light-weighted ages of J225546 appear to be slightly higher than the mass-weighted ages outside their error bounds, it is important to reiterate that one of the major caveats of the fossil-record method is that the stellar ages are indistinguishable for extremely old stellar populations. Furthermore, since the chosen isochrones of the SSP models used in the spectral fitting include ages up to Gyr, pPXF can estimate mass fractions at ages older than the Hubble time, depending on the assumed cosmology. Due to these limitations, we exercise caution in directly interpreting the individual absolute stellar ages and the star formation timescale. From the mean stellar ages, it is only clear that the values are low, and therefore suggest populations that are similar in age. However, the quantity is more related to intrinsic star formation and quenching in the galaxy, and less sensitive to dry mergers where stellar material is simply accreted without a subsequent star forming episode, unless the accreted galaxy itself is young. Therefore, the low values of the outer component do not necessarily confirm that the population was formed at the same time throughout the component, but could be attributed to stellar accretion from satellite galaxies that are approximately as old as the main galaxy.
@@
The mean mass-weighted stellar metallicities of the galaxies indicate metal-rich stellar populations, with super-solar metal abundances ([M/H]MW¿0). For the inner components, the metallicities are quite high compared to the global values for the galaxies. In contrast, the outer components signify metal-poor stellar populations, with sub-solar metal abundances ([M/H]MW¡0). Compared to the outer components, the inner components are significantly more metal-enhanced by factors of dex. Similarly, the light-weighted metallicities decrease inside-out, with the outer components showing less metal-enhanced stellar populations than their corresponding inner components by factors of dex. The relative differences between the mass-weighted and the light-weighted metallicities are quite small, between 0 and 0.2 dex, with the highest differences exhibited by the outer components. Our stellar populations analysis indicates a lack of recent star formation and younger stars, resulting in the mass and light-weighted maps appearing quite similar, with no concentrations of recent star formation. Any differences between the mass and light-weighted quantities are likely due to variations in metallicities rather than ages. These results align with the two-phase formation scenario where the inner component has likely assembled most of its stellar mass through in-situ bulk star formation when the interstellar medium was already enriched with metals, either through dissipative collapse of a gas cloud or through major mergers between gas-rich discs. This component would then accrete its stellar mass later in dry mergers of relatively metal-poor galaxies, such as dwarf galaxies, or through star formation resulting from relatively pristine gas inflow that dilutes the metal-enhanced interstellar medium.
@@
The consistency check for the stellar populations measurements using Lick indices are shown in Figure 5. The overlaid grid in grey represents the model SSP measurements from Vazdekis et al. (2010), and are conventionally used to convert the index measurements into physical stellar population properties. We have not done the latter conversion step, since most of the points lie outside the grid, instead only observing the qualitative implications of these results in tandem with the full-spectral fitting analysis. Nevertheless, the grid provides a guide to a ballpark estimate of the stellar age and metallicity. A cursory inspection of the points in the H - [MgFe]′ plane and their locations within the model SSP grid reveals that the inner components are typically older and more metal-rich than the outer components. However, the error bars are quite significant to conclusively determine this systematic difference, or if the inner components within each galaxy are older and more metal-enhanced compared to their respective outer components. The approximate stellar populations from this analysis align with the conclusions of those derived from the more extensive full-spectral fitting method.
@@
At this point, it is important to consider the possible caveats of employing the limited optical wavelength range of MUSE in deriving stellar populations. Given that the spectral range of the instrument extends only to the H absorption line in the blue end, this could lead to the implication that the lack of higher-order Balmer lines makes the analysis less sensitive to very recent star formation episodes at timescales Myr. We note that while the H and H would be indispensable in breaking the age-metallicity degeneracy when only the absorption line strengths are used in deriving stellar populations, this effect is mitigated in full-spectral fitting where multiple spectral features sensitive to both age and metallicity are modelled simultaneously alongside the stellar continuum. For this reason, and the fact that several galaxy components fall outside the model SSP grid in the H - [MgFe]′ plane, we conduct the rest of the stellar population analysis using the full-spectral fitting method.
@@ 4.3 Comparison with spatially resolved 2D stellar population maps
We also conducted a complementary stellar population analysis to compare the global properties described in Sect. 4.2. For this, the Voronoi binning technique was used to bin the IFU datacubes based on their S/N, and the spectra in each bin were summed together to retrieve a single spectrum with adequate S/N (we refer the reader to Sect. 3.4 for details). The binned spectra were then fit with pPXF using the same method detailed in Sect. 3.2. This technique allows us to create a 2D map of the stellar population properties and observe qualitatively any radial trends present in the galaxy. However, as stated in Sect. 3.4, these do not serve as a direct comparison to the results obtained with BUDDI due to the significant chance of mixing between different structural components. Nevertheless, these 2D maps offer an alternative way of comparing our results, albeit in a more generic way, and resemble more traditional analyses in the literature. The columns in Figure 6 represent the three elliptical galaxies, while each row shows a single stellar population property: mass-weighted age, light-weighted age, mass-weighted metallicity, and light-weighted metallicity in that order. The properties are mapped based on the colour bar above each plot, with the youngest ages and lowest metallicities in indigo, and the oldest ages and the highest metallicities in yellow. The black contours mark the 1 and 2 ellipses for the inner (dashed) and outer (solid) components modelled in the two-component fit to each galaxy.
@@
For galaxy J020536, the mass-weighted stellar age map shows a mildly negative gradient from the central region, corresponding to of the inner component derived from BUDDI, to the outskirts, corresponding to of the outer component. We find the negative mass-weighted age gradients to be composed of a dominant old ( Gyr) stellar population in the inner component, and a slightly younger but homogeneous population ( Gyr) on the outskirts. The light-weighted stellar age map however, does not appear to show any significant trends, similar to the findings of Lassen et al. (2021), where the spatially resolved light-weighted mean ages were derived through stellar population synthesis at the spaxel resolution. The mass-weighted and light-weighted stellar metallicity maps, on the other hand, show a very clear gradient from the centre to the outskirts. The core hosts the most metal-rich population in the galaxy, which becomes less enhanced moving radially outward towards the galactic outskirts. This trend is reflected in the inner component derived from BUDDI, with a negative gradient progressing from 1 to 2. This trend continues even moving toward 1 of the outer component, which hosts the least metal-enhanced populations in the galaxy. While both the mass-weighted and the light-weighted metallicity maps systematically show the same trend, the light-weighted map also appears to show less metal-enhanced populations compared to the mass-weighted map.
@@
Likewise, J205050 appears to host a central core that is older than the rest of the galaxy, which also lies within 1. Beyond the central region, however, the mass-weighted stellar age map is mostly homogeneous. From the light-weighted age map, it is difficult to observe any gradient, there appear to be both younger and older stellar populations across the outer component of the galaxy, but the oldest populations are still clustered in the centre. Both the mass-weighted and light-weighted metallicity maps effectively mirror those presented in galaxy J020536. The central core shows relatively higher metallicities, with a decline moving radially outward, again reflected in the modelled components moving from 1 to 2 of the inner component. The trends exhibited in the light-weighted metallicity map are again only different in the sense that they are systematically less metal-enhanced than in the mass-weighted map.
@@
The mass-weighted age map of J225546 does not indicate any substantial systematic trends, exhibiting both older stellar populations ( Gyr) and slightly younger populations with differences hardly reaching 1 Gyr across different regions of the galaxy. The light-weighted age map similarly shows a diverse range of stellar ages across the galaxy. In both cases, the region south-west of the galaxy shows an anomalous population compared to the rest of the galaxy outskirts or the outer component. This area corresponds to the location of the galaxy at , where the Voronoi bins are dominated by its strong emission rather than J225546, causing this region to appear younger. The metallicity maps are consistent with the previously described galaxies, with the central region dominated by metal-rich stars, and a negative gradient with lower metallicity stars being hosted in the outskirts. Similarly, the most metal-enhanced regions in both the mass-weighted and light-weighted maps are hosted within 2 of the modelled inner component; although the negative gradient is steeper in the light-weighted map than it is in the mass-weighted map moving from 1 to 2. We also note that the neighbouring galaxy hosts relatively metal-poor stellar populations; however analysing this galaxy is beyond the scope of this work.
@@ 4.4 Temporal mass and luminosity contributions from inner and outer components
The full spectral fitting process with pPXF returns a set of weights representing the relative contribution of each input SSP template to the optimised best-fit solution. From this, the 2D star formation history (SFH) of the galaxy components can be plotted on an age-[M/H] grid created from the stellar age and metallicity steps of the input stellar templates (see first and second columns in Figs. 7, 8, and 9). This proxy of SFH portrays the chemical evolution of the components through time, indicated by episodes of star formation. The 1D star formation histories (see third column in Figs. 7, 8, and 9) can then be constructed by cumulatively summing up the normalised weights over all the metallicities at each stellar age step in the grid, such that the total fraction is 1. In this study, since we analyse both mass-weighted and light-weighted stellar populations (Sect. 4.2), our reconstructed SFHs highlight the mass fraction and light fraction formed at different lookback times. The time evolution of the cumulative mass fraction provides insights into the stellar mass assembly history of the galaxy, such as whether it formed stars rapidly in the early universe or experienced prolonged star formation. The cumulative light fraction, on the other hand, highlights any recent star formation prominently since massive, young stars dominate the stellar luminosity of the galaxy and therefore have a higher impact on the light distribution when compared to the cumulative mass fraction. Furthermore, the light-weighted age distributions trace the most active phase of star formation in the galaxy.
@@
Figures 7, 8, and 9 illustrate the 1D (right-most panels) and 2D (first and middle panels) star formation histories from the mass-weighted and the light-weighted stellar populations in the top and bottom rows, respectively. The 2D SFHs are depicted by the mass or light fractions in the age-[M/H] grid, with the weights coloured according to the colour bar on the right of each plot. In the 1D SFHs, the inner component is shown in red and the outer component in blue. The axis in both depictions is time, but for the 2D SFH, the time is denoted as the logarithmic stellar age in yr, while for the 1D SFH, this is instead denoted by the lookback time in Gyr. For galaxy J225546, there is an additional “component” coloured in grey, which is in fact the neighbouring galaxy described in Sect. 3.4. This is only depicted here to emphasise the reliability of BUDDI in modelling multiple inherent components and external galaxies simultaneously, and it does not play a part in the inferences from the study.
@@ 4.4.1 Mass fractions
The weights from pPXF measure the fractions of stellar populations that contribute to the total stellar mass of the galaxy or object of interest. As a preliminary overview, in Table 8, we compute the mass contributions from stars that are older than 8 Gyr (), of intermediate ages between 4 and 8 Gyr (, and from stars that are younger than 3 Gyr (. These values are listed for the galaxy overall, and for the inner and outer components. Clearly, the majority of the mass of the ellipticals lies in stars older than 8 Gyr, in each component as well as collectively, making up over 85% of the mass for all except the outer component of J205050, which only contributes to 64%. The inner component appears to be entirely composed of such old stars, adding up to 100% of its mass. The fraction of stars younger than 3 Gyr is null or insignificant, except for the outer component of J225546, where they contribute mildly to 1.5% of the mass. This small fraction cannot be entirely ruled out as a consequence of the neighbouring object, which might not be completely deblended even when its profile is fit simultaneously with that of the target galaxy. The stars of intermediate age are prominent only in the outer component, varying from low to substantial in their mass contributions, between %. A more detailed analysis on the mass assembly of the components through cosmic time is described in the following Section.

### v2

@@ 4.2 Mean stellar populations
Figure 4 shows the stellar populations obtained from full spectral fitting with pPXF (see Sect. 3.2). The mean mass-weighted stellar metallicities [M/H] and the logarithmic stellar ages, are depicted in the upper panel, while the mean light-weighted properties are shown in the lower panel. The circles, squares, and triangles represent the three elliptical galaxies in the sample (J020536, J205050, and J225546, respectively), and the red and blue colours mark the inner and outer components derived from BUDDI in the two-component (Sérsic + Sérsic) models. The inner and outer components of the same galaxy are connected by the black lines for better readability. For comparison, the stellar population properties from the standard single component (Sérsic) model are shown in purple. Mass-weighted ages reflect the period when the majority of stellar mass was assembled and are less sensitive to the most luminous and younger stars; whereas light-weighted ages are influenced by recent star formation, which significantly impacts luminosity but contributes less to the overall mass (Hopkins, 2018).
@@
The mean mass and light-weighted ages of the galaxies (single Sérsic) and their individual components (inner and outer components in the double Sérsic model) are all older than 9 Gyr, which typically places them in the category of long-term quiescent ‘classical ellipticals’. The difference between the ages of the populations present in the inner and outer components ranges roughly between 1 and 5 Gyr. Plauchu-Frayn et al. (2012) define the star formation timescale as the difference between the mass-weighted and light-weighted stellar ages: Gyr. is high for objects that are continuously forming stars, while it is low for those that have ceased forming stars a long time ago. Following their definition, we compute for each component, as well as globally for each galaxy. Considering the ages of the galaxies as a whole, we find to be negligible, of the order of a few 100 Myr, well within their uncertainties. This insignificant difference is typically expected for classical ellipticals when considered as a global property, implying that their stellar populations are homogeneous in age and would have formed and quenched at roughly similar early look-back times. The inner and outer components similarly show virtually imperceptible values. While we recognise that the mean light-weighted ages of J225546 appear to be slightly higher than the mass-weighted ages outside their error bounds, it is important to reiterate that one of the major caveats of the fossil-record method is that the stellar ages are indistinguishable for extremely old stellar populations. Furthermore, since the chosen isochrones of the SSP models used in the spectral fitting include ages up to Gyr, pPXF can estimate mass fractions at ages older than the Hubble time, depending on the assumed cosmology. Due to these limitations, we exercise caution in directly interpreting the individual absolute stellar ages and the star formation timescale. From the mean stellar ages, it is only clear that the values are low, and therefore suggest populations that are similar in age. However, the quantity is more related to intrinsic star formation and quenching in the galaxy, and less sensitive to dry mergers where stellar material is simply accreted without a subsequent star forming episode, unless the accreted galaxy itself is young. Therefore, the low values of the outer component do not necessarily confirm that the population was formed at the same time throughout the component, but could be attributed to stellar accretion from satellite galaxies that are approximately as old as the main galaxy.
@@
The mean mass-weighted stellar metallicities of the galaxies indicate metal-rich stellar populations, with super-solar metal abundances ([M/H]MW¿0). For the inner components, the metallicities are quite high compared to the global values for the galaxies. In contrast, the outer components signify metal-poor stellar populations, with sub-solar metal abundances ([M/H]MW¡0). Compared to the outer components, the inner components are significantly more metal-enhanced by factors of dex. Similarly, the light-weighted metallicities decrease inside-out, with the outer components showing less metal-enhanced stellar populations than their corresponding inner components by factors of dex. The relative differences between the mass-weighted and the light-weighted metallicities are quite small, between 0 and 0.2 dex, with the highest differences exhibited by the outer components. Our stellar populations analysis indicates a lack of recent star formation and younger stars, resulting in the mass and light-weighted maps appearing quite similar, with no concentrations of recent star formation. Any differences between the mass and light-weighted quantities are likely due to variations in metallicities rather than ages. These results align with the two-phase formation scenario where the inner component has likely assembled most of its stellar mass through in situ bulk star formation when the interstellar medium was already enriched with metals, either through dissipative collapse of a gas cloud or through major mergers between gas-rich discs. This component would then accrete its stellar mass later in dry mergers of relatively metal-poor galaxies, such as dwarf galaxies, or through star formation resulting from relatively pristine gas inflow that dilutes the metal-enhanced interstellar medium.
@@
The consistency check for the stellar populations measurements using Lick indices are shown in Fig. 5. The overlaid grid in grey represents the model SSP measurements from Vazdekis et al. (2010), and are conventionally used to convert the index measurements into physical stellar population properties. We have not done the latter conversion step, since most of the points lie outside the grid, instead only observing the qualitative implications of these results in tandem with the full-spectral fitting analysis. Nevertheless, the grid provides a guide to a ballpark estimate of the stellar age and metallicity. A cursory inspection of the points in the H - [MgFe]′ plane and their locations within the model SSP grid reveals that the inner components are typically older and more metal-rich than the outer components. However, the error bars are quite significant to conclusively determine this systematic difference, or if the inner components within each galaxy are older and more metal-enhanced compared to their respective outer components. The approximate stellar populations from this analysis align with the conclusions of those derived from the more extensive full-spectral fitting method.
@@
At this point, it is important to consider the possible caveats of employing the limited optical wavelength range of MUSE in deriving stellar populations. Given that the spectral range of the instrument extends only to the H absorption line in the blue end, this could lead to the implication that the lack of higher-order Balmer lines makes the analysis less sensitive to very recent star formation episodes at timescales Myr. We note that while the H and H would be indispensable in breaking the age-metallicity degeneracy when only the absorption line strengths are used in deriving stellar populations, this effect is mitigated in full-spectral fitting where multiple spectral features sensitive to both age and metallicity are modelled simultaneously alongside the stellar continuum. For this reason, and the fact that several galaxy components fall outside the model SSP grid in the H - [MgFe]′ plane, we conducted the rest of the stellar population analysis using the full-spectral fitting method.
@@ 4.3 Comparison with spatially resolved 2D stellar population maps
We also conducted a complementary stellar population analysis to compare the global properties described in Sect. 4.2. For this, the Voronoi binning technique was used to bin the IFU datacubes based on their S/N, and the spectra in each bin were summed together to retrieve a single spectrum with adequate S/N (we refer the reader to Sect. 3.4 for details). The binned spectra were then fitted with pPXF using the same method detailed in Sect. 3.2. This technique allows us to create a 2D map of the stellar population properties and observe qualitatively any radial trends present in the galaxy. However, as is stated in Sect. 3.4, these do not serve as a direct comparison to the results obtained with BUDDI due to the significant chance of mixing between different structural components. Nevertheless, these 2D maps offer an alternative way of comparing our results, albeit in a more generic way, and resemble more traditional analyses in the literature. The columns in Fig. 6 represent the three elliptical galaxies, while each row shows a single stellar population property: mass-weighted age, light-weighted age, mass-weighted metallicity, and light-weighted metallicity, in that order. The properties are mapped based on the colour bar above each plot, with the youngest ages and lowest metallicities in indigo, and the oldest ages and the highest metallicities in yellow. The black contours mark the 1 and 2 ellipses for the inner (dashed) and outer (solid) components modelled in the two-component fit to each galaxy.
@@
For galaxy J020536, the mass-weighted stellar age map shows a mildly negative gradient from the central region, corresponding to of the inner component derived from BUDDI, to the outskirts, corresponding to of the outer component. We find the negative mass-weighted age gradients to be composed of a dominant old ( Gyr) stellar population in the inner component, and a slightly younger but homogeneous population ( Gyr) on the outskirts. The light-weighted stellar age map, however, does not appear to show any significant trends, similar to the findings of Lassen et al. (2021), where the spatially resolved light-weighted mean ages were derived through SPS at the spaxel resolution. The mass-weighted and light-weighted stellar metallicity maps, on the other hand, show a very clear gradient from the centre to the outskirts. The core hosts the most metal-rich population in the galaxy, which becomes less enhanced moving radially outwards towards the galactic outskirts. This trend is reflected in the inner component derived from BUDDI, with a negative gradient progressing from 1 to 2. This trend continues even moving towards 1 of the outer component, which hosts the least metal-enhanced populations in the galaxy. While both the mass-weighted and the light-weighted metallicity maps systematically show the same trend, the light-weighted map also appears to show less metal-enhanced populations compared to the mass-weighted map.
@@
Likewise, J205050 appears to host a central core that is older than the rest of the galaxy, which also lies within 1. Beyond the central region, however, the mass-weighted stellar age map is mostly homogeneous. From the light-weighted age map, it is difficult to observe any gradient, there appear to be both younger and older stellar populations across the outer component of the galaxy, but the oldest populations are still clustered in the centre. Both the mass-weighted and light-weighted metallicity maps effectively mirror those presented in galaxy J020536. The central core shows relatively higher metallicities, with a decline moving radially outwards, again reflected in the modelled components moving from 1 to 2 of the inner component. The trends exhibited in the light-weighted metallicity map are again only different in the sense that they are systematically less metal-enhanced than in the mass-weighted map.
@@
The mass-weighted age map of J225546 does not indicate any substantial systematic trends, exhibiting both older stellar populations ( Gyr) and slightly younger populations with differences hardly reaching 1 Gyr across different regions of the galaxy. The light-weighted age map similarly shows a diverse range of stellar ages across the galaxy. In both cases, the region south-west of the galaxy shows an anomalous population compared to the rest of the galaxy outskirts or the outer component. This area corresponds to the location of the galaxy at , where the Voronoi bins are dominated by its strong emission rather than J225546, causing this region to appear younger. The metallicity maps are consistent with the previously described galaxies, with the central region dominated by metal-rich stars, and a negative gradient with lower metallicity stars being hosted in the outskirts. Similarly, the most metal-enhanced regions in both the mass-weighted and light-weighted maps are hosted within 2 of the modelled inner component; although the negative gradient is steeper in the light-weighted map than it is in the mass-weighted map moving from 1 to 2. We also note that the neighbouring galaxy hosts relatively metal-poor stellar populations; however, analysing this galaxy is beyond the scope of this work.
@@ 4.4 Temporal mass and luminosity contributions from inner and outer components
The full spectral fitting process with pPXF returns a set of weights representing the relative contribution of each input SSP template to the optimised best-fit solution. From this, the 2D SFH of the galaxy components can be plotted on an age-[M/H] grid created from the stellar age and metallicity steps of the input stellar templates (see first and second columns in Figs. 7, 8, and 9). This proxy of SFH portrays the chemical evolution of the components through time, indicated by episodes of star formation. The 1D SFHs (see third column in Figs. 7, 8, and 9) can then be constructed by cumulatively summing up the normalised weights over all the metallicities at each stellar age step in the grid, such that the total fraction is 1. In this study, since we analyse both mass-weighted and light-weighted stellar populations (Sect. 4.2), our reconstructed SFHs highlight the mass fraction and light fraction formed at different look-back times. The time evolution of the cumulative mass fraction provides insights into the stellar mass assembly history of the galaxy, such as whether it formed stars rapidly in the early universe or experienced prolonged star formation. The cumulative light fraction, on the other hand, highlights any recent star formation prominently since massive, young stars dominate the stellar luminosity of the galaxy and therefore have a higher impact on the light distribution when compared to the cumulative mass fraction. Furthermore, the light-weighted age distributions trace the most active phase of star formation in the galaxy.
@@
Figures 7, 8, and 9 illustrate the 1D (right-most panels) and 2D (first and middle panels) SFHs from the mass-weighted and the light-weighted stellar populations in the top and bottom rows, respectively. The 2D SFHs are depicted by the mass or light fractions in the age-[M/H] grid, with the weights coloured according to the colour bar on the right of each plot. In the 1D SFHs, the inner component is shown in red and the outer component in blue. The axis in both depictions is time, but for the 2D SFH, the time is denoted as the logarithmic stellar age in years, while for the 1D SFH, this is instead denoted by the look-back time in gigayears. For galaxy J225546, there is an additional ‘component’ coloured in grey, which is in fact the neighbouring galaxy described in Sect. 3.4. This is only depicted here to emphasise the reliability of BUDDI in modelling multiple inherent components and external galaxies simultaneously, and it does not play a part in the inferences from the study.
@@ 4.4.1 Mass fractions
The weights from pPXF measure the fractions of stellar populations that contribute to the total stellar mass of the galaxy or object of interest. As a preliminary overview, in Table 8, we compute the mass contributions from stars that are older than 8 Gyr (), of intermediate ages between 4 and 8 Gyr (, and from stars that are younger than 3 Gyr (. These values are listed for the galaxy overall, and for the inner and outer components. Clearly, the majority of the mass of the ellipticals lies in stars older than 8 Gyr, in each component as well as collectively, making up over 85% of the mass for all except the outer component of J205050, which only contributes to 64%. The inner component appears to be entirely composed of such old stars, adding up to 100% of its mass. The fraction of stars younger than 3 Gyr is null or insignificant, except for the outer component of J225546, where they contribute mildly to 1.5% of the mass. This small fraction cannot be entirely ruled out as a consequence of the neighbouring object, which might not be completely deblended even when its profile is fitted simultaneously with that of the target galaxy. The stars of intermediate age are prominent only in the outer component, varying from low to substantial in their mass contributions, between %. A more detailed analysis on the mass assembly of the components through cosmic time is described in the following section.